OpenStudy (anonymous):

How many solutions exist for the system, x + y = 1 and y = -x + 1? no solution one solution infinitely many solutions

OpenStudy (anonymous):

Solve the first equation for y; that is, put it in slope intercept form.

OpenStudy (anonymous):

x+y=1 Since x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting x from both sides. y=-x+1 To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept. y=mx+b Using the y=mx+b formula, m=-1 and b=1. m=-1, b=1

OpenStudy (anonymous):

x+y=1 Since x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting x from both sides. y=-x+1